* Return to course ÈBusiness Course Question 7 Not yet answered Marked out of 5.00 P Flag question 2020 Performing a Vertical please show work
Answer #1

Vertifical Analysis

2020 2019
Income Statement 2020 2019 % of Sales % of Sales
Revenues $            39,000 $            42,900 100.0% 100.0%
Cost of Goods Sold $            25,920 $            29,160 66.5% 68.0%
Gross Margin $            13,080 $            13,740 33.5% 32.0%
Operating Expenses $               8,800 $            10,000 22.6% 23.3%
Net Income $              4,280 $              3,740 11.0% 8.7%

Calculations

2020 2019
% of Sales % of Sales
=39000/39000 =42900/42900
=25920/39000 =29160/42900
=13080/39000 =13740/42900
=8800/39000 =10000/42900
=4280/39000 =3740/42900

Horizontal Analysis

Dollar Percentage
Income Statement 2020 2019 Change Change
Revenues $            39,000 $            42,900 $             (3,900) -9.1%
Cost of Goods Sold $            25,920 $            29,160 $             (3,240) -11.1%
Gross Margin $            13,080 $            13,740 $                (660) -4.8%
Operating Expenses $               8,800 $            10,000 $             (1,200) -12.0%
Net Income $              4,280 $              3,740 $                  540 14.4%

Calculations

Dollar Percentage
Change Change
=39000-42900 =-3900/42900
=25920-29160 =-3240/29160
=13080-13740 =-660/13740
=8800-10000 =-1200/10000
=4280-3740 =540/3740

If you have any query, kindly comment with your query and please mark thumbs up.